# Reconstructing Intelligible Speech from the Pressure Sensor Data in HVACs

### v1

@@ 5 Threat Model
a) Attacker’s goal: The attacker can eavesdrop on the conversation with unrestricted vocabulary using HVAC systems where DPSs are located close to human occupancy. In addition to eavesdropping a natural conversation, the attacker can also understand how many people are present inside of a room/facility, how many of them are males and females, what types of sound sources are present, and from the sound sources, the attacker can also identify what types of activities are going on inside the room/facility. The attacker collects pressure data from DPSs and reconstructs intelligible audio from the noisy and aliased pressure data using our WaLi. WaLi uses complex-valued models to reconstruct both magnitude and phase to generate intelligible audio for at least 0.5 kHz sampling frequency.
@@
b) Attacker’s access level: Prior works with the capability to spy on natural conversations use malicious apps to access inertial sensors [10, 13] to collect data for eavesdropping. These attacks require the installation of such apps to initiate the eavesdropping. In contrast, our attack model exploits the normal behavior of HVACs without installing additional software on HVACs. Access to collect pressure data from DPSs is possible in the following scenarios.
@@
First, an attacker disguised as a malicious employee or maintenance person can access pressure data from the Building Management System (BMS) software dashboard, as in modern buildings, pressure sensors are integrated into the BMS using standard protocols, such as Modbus TCP, KNX, and LonWorks.
@@
Second, in many cases, the BMS is handled by a third-party vendor, HVAC contractors, or system integrator, especially in commercial buildings, hospitals, labs, and large campuses. These third-party vendors manage complex integration, security and control, compliance, and 24/7 monitoring. In many cases, authorities often outsource teams to provide continuous support and alert handling. An attacker disguised as one of these third-party vendors can access sensitive pressure data via a web-based interface, historical logs, or an Open Platform Communication (OPC) server.
@@
Fourth is a potential supply chain attack, which has been rumored to occur in the past [47, 48, 49, 50] and was recently demonstrated to be feasible [51], involving tampering with HVAC systems during delivery or installation. In such an attack, a competitor could intercept the equipment and modify DPSs by embedding a data collection device with a radio transmitter inside, and then allow the delivery or installation to proceed as planned to the commercial facility.
c) Non-invasive attack: Our eavesdropping attack using the HVAC system is non-invasive as it is performed without making direct physical contact with the target DPS. The attacker doesn’t need to directly open or physically touch DPSs. However, we expect that attackers can examine the behavior of a similar sensor subjected to acoustic impacts before initiating an actual attack.
d) Attacker’s resources and cost: We assume that the attacker knows how the HVAC system works and how the data can be collected from pressure sensors.
e) Assumption: Considering the practicality of the attack, we assume that the attacker has sufficient pressure sensor data collected from the victim and does not have matching ground truth audio from the victim. WaLi can reconstruct the audio even if the speaker is different. This makes our attack model flexible compared to prior works [13, 34], where prior works assume that clean ground-truth speech is available via a microphone or a phishing call.
Figure 6: WaLi has complex-valued encoders, decoders, complex-valued skip blocks, CGAB, and complex multireoslution STFT loss.
6 WaLi ARCHITECTURE DESIGN
@@
a) WaLi processes the input pressure sensor data in the complex-valued T-F spectrum and enhances both the amplitude and phase of the pressure sensor data to reconstruct intelligible speech from the aliased pressure sensor data.
b) WaLi handles transient interference in noisy conditions of the pressure sensor data using joint magnitude and phase reconstruction using a complex-valued network.
c) WaLi reconstructs speech from the victim’s pressure sensor data without using the ground truth audio of the victim.
WaLi adopts a complex-valued U-Net model and processes the incoming low-resolution and noisy pressure data using the complex-valued T-F spectrogram. We will first explain the reasons behind the use of the complex-valued T-F spectrogram and network.
Why the complex-valued spectrogram and network: A complex-valued spectrogram is a T-F representation of a signal that contains both magnitude and phase information, not just the energy or amplitude as in a standard real-valued spectrogram. A complex-valued T-F spectrogram is typically generated by the Short-Time Fourier Transform (STFT) and then outputs a complex value $S(t,f)$ at each time-frequency bin, following Eq. 1.
@@
where $A(t,f)$ is the magnitude and $\phi(t,f)$ is the phase of the spectrum. Prior works [13, 10] only use real-valued T-F spectrograms. Therefore, they cannot predict phase information and hence, need a vocoder to generate audio from real-valued spectrograms. Moreover, real-valued T-F spectrogram-based methods typically use Mean Square Error (MSE) loss for magnitudes and cannot use phase-reconstruction loss functions to improve the phase quality while reconstructing speech from the low-resolution pressure sensor data. Therefore, motivated by the fact that phase plays a crucial role in speech enhancement [12], WaLi adopts a joint reconstruction of magnitudes and phases. As complex-valued spectrograms have both magnitude and phase information (refer to Eqn. 1), WaLi takes complex-valued T-F spectrograms as input and generates complex-valued T-F spectrograms at its output. To process the complex-valued T-F spectrograms, we design WaLi as a complex-valued network having U-Net as its backbone.
WaLi architecture: The detailed architecture of the proposed WaLi is shown in Fig. 6. The network consists of four main components: (i) a total of 16 full complex-valued encoder-decoder blocks (i.e., 8 encoders and 8 decoders), (ii) complex-valued skip blocks, (iii) complex-valued conformer in the bottleneck layer, and (iv) complex-valued global attention blocks - CGAB. The complex domain processing by our proposed WaLi has the potential to recover clean phases from the noisy pressure sensor and to reconstruct magnitudes from the aliased spectrum (i.e., pressure sensor is sampled at a minimum of 0.5 kHz).
6.1 Complex-Valued Encoders
The complex-valued spectrogram of the pressure sensor data is transformed to real and imaginary components, and then the real and imaginary components are combined into a complex-valued tensor, which serves as input to the first complex encoder block (E1). Formally, the input low-resolution pressure data $R_{in}$ is first transformed into STFT spectrogram, denoted by $S_{in}$ in Fig. 6. Here, $S_{in}(=S^{r}+jS^{i})\in\mathbb{C}^{F\times T}$ is a complex-valued spectrogram, where $F$ denotes the number of frequency bins and $T$ denotes the number of time frames.
Typically, each encoder in a U-Net extracts hierarchical features from the input or previous layers by downsampling the features using a convolution operation. WaLi adopts complex-valued encoders, which is built upon complex-valued convolution to ensure successive extraction of both magnitude and phase from the complex-valued T-F spectrogram. Complex convolution is the key difference between a complex-valued encoder and a real-valued encoder. Formally, the complex-valued T-F spectrogram $S_{in}$ is fed into 2D complex convolution layers of the first encoder to produce feature $S_{1}\in\mathbb{C}^{F\times T\times C}$, where C is the number of channels. If a complex kernel is denoted by $W={W}_{r}+j{W}_{i}$, the complex convolution is defined by Eqn. 2.
@@ 6.1 Complex-Valued Encoders
where * denotes the convolution, $S^{r}_{1}$\&& $S^{i}_{1}$ are real and imaginary parts of $S_{1}$, and ${b}_{r}$\&& ${b}_{i}$ are bias terms.
The convolution output is then normalized using Complex Batch Normalization (CBN) for stable training. CBN is an extension of traditional batch normalization to complex-valued neural networks, where inputs, weights, or features are complex numbers (i.e., having both real and imaginary parts). Instead of treating real and imaginary parts independently, CBN treats them as a 2D vector and normalizes using a 2×2 covariance matrix using Eqn. 3.
@@
where $\Sigma_{S_{1}}^{-\frac{1}{2}}$ is the inverse square root matrix of $\Sigma_{S_{1}}$ used to decorrelate and normalize the input $S_{1}$, and $\mu$ is the complex mean vector of $S_{1}$.
Next, the output from CBN passes through a complex ReLU activation for adding nonlinearity. Formally, first encoder output, denoted by $E_{1}$ can be expressed by Eqn. 5.
@@
The output of the first encoder $E_{1}$ is given as input to the 2nd encoder, and 2nd encoder’s output to the third encoder, and so on. Every complex encoder has a similar complex 2D convolution (Eqn. 2), CBN (Eqn. 4), and complex ReLU activation (Eqn. 5).
6.2 Complex-Valued Conformer in Bottlenecks
We use complex-valued conformers in the bottleneck layer of our WaLi to capture local and global dependencies among consecutive spectrograms. A conformer [9] combines the strengths of CNNs and Transformers. Traditional Transformers are great at capturing long-term dependencies, but struggle with local patterns that exist in each phoneme. Convolutional layers, on the other hand, excel at local context but lack long-range modeling. The conformer blends both by stacking self-attention and convolutional modules in each layer, resulting in a highly expressive and efficient model for sequential input. Our complex-valued conformer optimally balances global context with fine-grained local information in phase and amplitude domains for the successful reconstruction of intelligible audio from pressure data.
The output of the last (eighth) encoder block $E_{8}$ is expressed as $E_{8}\in\mathbb{R}^{B\times C\times F\times T}$ and fed as input into the conformer block. Here, $B$ denotes the batch size, $C$ the number of channels, $F$ the spectral dimension, and $T$ the temporal dimension. Next, we rearrange $E_{8}$ using a permutation and reshaping so that the conformer processes each channel independently. For a given channel, the complex-valued conformer comprises complex multi-head self-attention (MHSA), complex feed-forward (FF), and complex convolutional modules, inspired by [9]. The complex MHSA uses complex-valued queries, keys, and values over the real and imaginary parts. The complex FF module uses complex linear layers, complex ReLU activation, and CBN.
6.3 Complex-Valued Decoder
The output from the bottleneck layer, denoted by $\mathbf{Z}$ (i.e., latent space), is given as input to the first decoder. The decoder reconstructs the T-F representation from the latent space by employing a complex transposed convolution by upsampling by a factor of two at each decoder block. For a latent complex tensor $\mathbf{Z}=\mathbf{Z}_{r}+j\,\mathbf{Z}_{i},$ the transposed convolution is formulated as:
@@ 6.3 Complex-Valued Decoder
where $\tilde{\mathbf{Y}}_{1}^{r}$\&& $\tilde{\mathbf{Y}}_{1}^{i}$ are real and imaginary parts of $\tilde{\mathbf{Y}}_{1}$ (i.e., output after the transposed convolution in the first decoder), and ${b}_{r}$\&& ${b}_{i}$ are bias terms. $\tilde{\mathbf{Y}}_{1}$ is then normalized by CBN and activated by a complex ReLU activation as Eqn. 7.
@@
where $D_{1}$ is the output of the first decoder, which is given as input to the 2nd encoder, and so on. Every complex decoder has a similar complex 2D convolution (Eqn. 6), CBN (Eqn. 4), and complex ReLU activation (Eqn. 7).
6.4 Complex Skip Block
A skip connection in our proposed WaLi passes high-dimensional features from complex-valued encoders to the appropriate decoders. This enables the model to preserve the spatial features, which may lost during the down-sampling operation, and guides the network to propagate from encoders to decoders. WaLi implements skip blocks in complex domains, inspired by [52], to enable the proper flow of complex features from the encoder’s output to decoders. Each complex skip block applies a complex convolution on the encoder output, followed by a CBN and a complex ReLU activation. Formally, the complex skip block’s output, denoted by $SK_{n}$ is (i.e, n is 1 to 8):
@@ 6.4 Complex Skip Block
6.5 Complex Global Attention Block (CGAB)
Long-range correlations exist along both the time and the frequency axes in a complex-valued T-F spectrogram. As the audio signal embedded in the pressure sensor data is a time series signal, inter-phoneme correlations exist along the time axis. Moreover, harmonic correlations also exist among pitch and formants along the frequency axis. As a convolution kernel is limited by its receptive field, standard convolutions cannot capture global correlations on the time and frequency axes in a complex-valued T-F spectrogram. Therefore, we propose the Complex Global Attention Block (CGAB) to capture long-range correlation from the T-F axes. Please note that Frequency Transformation Blocks (FTBs) used in [10] published in USENIX, 2024, don’t work along both the T-F axes.
The detailed implementation of our proposed CGAB is shown in Fig. 7. CGAB provides attention to the time and frequency axes of a complex-valued T-F spectrogram by following two steps:
Step 1 - Reshaping along the T-F axes: The output $E_{n}$ from the encoder is decomposed in 2 steps by CGAB into two tensors: one along the time axis and another along the frequency axis. Formally, $E_{n}$, which has a feature dimension of $C\times F\times T$, is given at the input of CGAB. At the first stage of reshaping, $E_{n}$ parallelly reshaped into $C.T$ vectors with dimension $C\cdot T\times F$ and into $C.F$ vectors with dimension $C\cdot F\times T$. This reshaping is done using 2D complex convolution, CBN, and complex ReLU activation, followed by vector reshaping. In the second stage of reshaping, $C\cdot T\times F$ is reshaped into $1\times T\times F$ and $C\cdot F\times T$ is reshaped into $1\times F\times T$ using 1D complex convolution, CBN, complex ReLU activation followed by vector reshaping. The tensors with dimension $1\times F\times T$ capture the global harmonic correlation along the frequency axis and $1\times T\times F$ capture the global inter-phoneme correlation along the time axis. The captured features along the T-F axes and the original features from $E_{n}$ are point-wise multiplied together to generate a combined feature map with a dimension of $C\times T\times F$ and $C\times F\times T$ along T and F axes, respectively. This point-wise multiplication captures the inter-channel relationship between the encoder’s output and complex time and frequency axes.
Step 2 - Global attention along the T-F axes: It is possible to treat the spectrogram as a 2D image and learn the correlations between every two pixels in the 2D image. However, this is computationally too costly and is not realistic. On the other hand, ideally, we can use self-attention [53] to learn the attention map from two consecutive complex T-F spectrograms. But this might not be necessary. Because, on the time axis in each T-F spectrogram, when calculating SNR, the same set of parameters in the recursive relation are used, which suggests that temporal correlation is time-invariant among consecutive spectrograms. Moreover, harmonic correlations are independent in the consecutive spectrograms [54].
@@ 6.5 Complex Global Attention Block (CGAB)
Based on this understanding, we propose a self-attention technique along the T-F axes within each spectrogram, without considering correlations among consecutive spectrograms. Specifically, attention on frequency and time axes are implemented by two separate fully connected (FC) layers. Along the time path, the input and output dimensions of FC layers are $C\times T\times F$. Along the frequency path, the input and output dimensions of FC layers are $C\times F\times T$. FC layer learns weights from complex T-F spectrograms and technically is different from the self-attention [53] operation. To capture interchannel relationships among the input $E_{n}$ and output of FC layers, concatenation happens, followed by 2D complex convolutions, CBN, and complex ReLU activation. Finally, the learned weights from the T-F axes are concatenated together to form a unified tensor, which holds joint information on the T-F global correlations from each spectrogram.
We use only two CGABs - one after the 1st encoder, and another after the 7th encoder.
6.6 Complex Multiresolution STFT Loss

### v2

@@ III Threat Model
b) Attacker’s goal: The attacker can eavesdrop on speech with unrestricted vocabulary using HVAC systems where DPSs are located close to human occupancy. In addition to eavesdropping a natural speech, the attacker may also understand how many people are present inside a room/facility, how many of them are males and females, what types of sound sources are present, and from the sound sources, the attacker can also identify what types of activities are going on inside.
@@
c) Attacker’s access level: Prior works require installation of malicious apps to access inertial sensors [10, 13] to collect data for eavesdropping. In contrast, our attack model exploits the normal behavior of HVACs without installing additional software on HVACs. Access to collect pressure data from DPSs is possible in the following scenarios.
@@
First, an attacker disguised as a malicious employee or maintenance person can access pressure data from the Building Management System (BMS) software dashboard, as in modern buildings, pressure sensors are integrated into the BMS using standard protocols, such as Modbus TCP, and KNX [25].
@@
Second, in many cases, the BMS is handled by third-party contractors, or system integrators, especially in commercial buildings, hospitals, labs, and large campuses. In many cases, authorities often outsource teams to provide continuous support and alert handling. An attacker disguised as one of these third-party vendors can access sensitive pressure data via a web-based interface, historical logs, or an Open Platform Communication (OPC) server.
@@
d) Non-invasive attack: Our eavesdropping attack using the HVAC is non-invasive as it is performed without making direct physical contact with the target DPS. However, we expect that attackers can examine the behavior of a similar sensor subjected to acoustic impacts before initiating an actual attack.
e) Attacker’s resources and cost: We assume that the attacker knows how the HVAC system works and how the data can be collected from pressure sensors.
f) Assumption: Considering the practicality of the attack, we assume that the attacker has sufficient pressure sensor data collected from the victim and does not have matching ground truth audio from the victim. WaLi can reconstruct the audio even if the speaker is different. This makes our attack model flexible compared to prior works [13, 24], where prior works assume that clean ground-truth speech is available via a microphone or a phishing call.
Figure 5: WaLi has complex-valued encoders, decoders, complex-valued skip blocks, CGAB, and complex multiresolution STFT loss.
IV WaLi Architecture Design
@@
a) WaLi processes the pressure sensor data in the complex-valued T-F spectrum and enhances both the amplitude and phase of the pressure data to reconstruct intelligible speech.
b) WaLi handles transient interference in pressure sensor data using joint magnitude and phase reconstruction using a complex-valued T-F spectrogram and network.
Why the complex-valued spectrogram and network: A complex-valued T-F spectrogram of a signal contains both magnitude and phase information, not just the energy or amplitude as in a standard real-valued spectrogram. A complex-valued T-F spectrogram is typically generated by the Short-Time Fourier Transform (STFT) and then outputs a complex value $S(t,f)$ at each time-frequency bin, following Eq. 1.
@@
where $A(t,f)$ is the magnitude and $\phi(t,f)$ is the phase of the spectrum. Prior works [13, 10] only use real-valued T-F spectrograms. Therefore, they cannot predict phase and hence, need a vocoder [26] to generate audio from real-valued spectrograms. Moreover, real-valued T-F spectrogram-based methods typically use Mean Squared Error (MSE) loss for magnitudes and cannot use phase-reconstruction loss functions [27] to improve the phase quality while reconstructing speech from the low-resolution pressure sensor data. Therefore, motivated by the fact that phase plays a crucial role in speech enhancement [12], WaLi adopts a joint reconstruction of magnitudes and phases. mycolorAs complex-valued spectrograms have both magnitude and phase information (refer to Eqn. 1), WaLi takes complex-valued T-F spectrograms as input and generates complex-valued T-F spectrograms at its output.
WaLi architecture: To process the complex-valued T-F spectrograms, we design WaLi as a complex-valued network having a U-Net at its backbone (see Fig. 5) with four main components: (i) 16 full complex-valued encoder-decoder blocks (i.e., 8 encoders and 8 decoders), (ii) complex-valued skip blocks, (iii) complex-valued conformer in the bottleneck layer, and (iv) complex-valued global attention blocks - CGAB. The complex domain processing by WaLi has the potential to recover clean phases from the noisy pressure sensor and to reconstruct magnitudes from the aliased spectrum (i.e., pressure is sampled at a minimum of 0.5 kHz).
IV-A Complex-Valued Encoders
The pressure sensor data is transformed to complex-valued T-F spectrogram by STFT and given as input to the first complex encoder block (E1). Formally, the input low-resolution pressure data $R_{in}$ is first transformed into STFT spectrogram, denoted by $S_{in}$ in Fig. 5. Here, $S_{in}(=S^{r}+jS^{i})\in\mathbb{C}^{F\times T}$ is a complex-valued spectrogram, where $F$ denotes the number of frequency bins and $T$ denotes the number of time frames.
WaLi adopts complex-valued encoders, which is built upon complex-valued convolution to ensure successive extraction of both magnitude and phase from the complex-valued T-F spectrogram. Complex convolution is the key difference between a complex-valued encoder and a real-valued encoder. Formally, the complex-valued T-F spectrogram $S_{in}$ is fed into 2D complex convolution layers of the first encoder to produce feature $S_{1}\in\mathbb{C}^{F\times T\times C}$, where C is the number of channels. If a complex kernel is denoted by $W={W}_{r}+j{W}_{i}$, the complex convolution is defined by Eqn. 2.
@@ IV-A Complex-Valued Encoders
where * denotes the convolution, $S^{r}_{1}$ &\& $S^{i}_{1}$ are real and imaginary parts of $S_{1}$, and ${b}_{r}$ &\& ${b}_{i}$ are bias terms.
The convolution output is then normalized using Complex Batch Normalization (CBN) for stable training. CBN is an extension of traditional batch normalization to complex-valued neural networks, where inputs, weights, or features have both real and imaginary parts. Instead of treating real and imaginary parts independently, CBN treats them as a 2D vector and normalizes using a 2×2 covariance matrix using Eqn. 3.
@@
where $\Sigma_{S_{1}}^{-\frac{1}{2}}$ is the inverse square root matrix of $\Sigma_{S_{1}}$ used to decorrelate and normalize the input $S_{1}$, and $\mu$ is the complex mean vector of $S_{1}$. Next, the output from CBN passes through a complex ReLU activation for adding nonlinearity. Formally, first encoder output, denoted by $E_{1}$ is expressed by Eqn. 5.
@@
The output of the 1st encoder is given as input to the 2nd encoder, and 2nd encoder’s output to third encoder, and so on. Each complex encoder has a similar complex 2D convolution (Eqn. 2), CBN (Eqn. 4), and complex ReLU (Eqn. 5).
IV-B Complex-Valued Conformer in Bottlenecks
We design complex-valued conformers in the bottleneck layer because conformers [9] combine the strengths of CNNs and Transformers by stacking self-attention and convolutional modules in each layer, resulting in a highly expressive and efficient model for sequential input. Our complex-valued conformer optimally balances global context with fine-grained local information in phase and amplitude domains for the successful reconstruction of audio from pressure data.
The output of the last (eighth) encoder block, expressed as $E_{8}\in\mathbb{R}^{B\times C\times F\times T}$, is fed as input into the conformer. Here, $B$ denotes the batch size, $C$ the number of channels, $F$ the spectral dimension, and $T$ the temporal dimension. We rearrange $E_{8}$ using a permutation and reshaping so that the conformer processes each channel independently. For a given channel, the complex-valued conformer comprises complex multi head self-attention (MHSA), complex feed-forward (FF), and complex convolutional modules. The complex MHSA uses complex-valued queries, keys, and values over the real and imaginary parts. The complex FF module uses complex linear layers, complex ReLU activation, and CBN.
IV-C Complex-Valued Decoder
The output from the bottleneck layer, denoted by $\mathbf{Z}$ (i.e., latent space), is given as input to the first decoder. The decoder reconstructs the T-F representation from the latent space by employing a complex transposed convolution (upsampling by a factor of two) at each decoder. For a latent complex tensor $\mathbf{Z}=\mathbf{Z}_{r}+j\,\mathbf{Z}_{i},$ the transposed convolution is formulated as:
@@ IV-C Complex-Valued Decoder
where $\tilde{\mathbf{Y}}_{1}^{r}$ &\& $\tilde{\mathbf{Y}}_{1}^{i}$ are real and imaginary parts of $\tilde{\mathbf{Y}}_{1}$ (i.e., output after the transposed convolution in the first decoder), and ${b}_{r}$ &\& ${b}_{i}$ are bias terms. $\tilde{\mathbf{Y}}_{1}$ is then normalized by CBN and activated by a complex ReLU activation as Eqn. 7.
@@
where $D_{1}$ is the output of the first decoder, which is given as input to the 2nd decoder, and so on. Every complex decoder has a similar complex 2D convolution (Eqn. 6), CBN (Eqn. 4), and complex ReLU activation (Eqn. 7).
IV-D Complex Skip Block
WaLi implements skip blocks in complex domains, inspired by [28], to enable the proper flow of complex-valued high-dimensional features from the encoder’s output to appropriate decoders. Each complex skip block applies a complex convolution on the encoder output, followed by a CBN and a complex ReLU activation. Formally, the complex skip block’s output, denoted by $SK_{n}$ is (i.e, n is 1 to 8):
@@ IV-D Complex Skip Block
IV-E Complex Global Attention Block (CGAB)
Long-range correlations exist along both the time and the frequency axes in a complex-valued T-F spectrogram. As the audio signal embedded in the pressure sensor data is a time series signal, inter-phoneme correlations exist along the time axis. Moreover, harmonic correlations also exist among pitch and formants along the frequency axis. As a convolution kernel is limited by its receptive field, standard convolutions cannot capture global correlations on the time and frequency axes in a complex-valued T-F spectrogram. Therefore, we propose a Complex Global Attention Block (CGAB) to capture long-range correlation from both the T-F axes. Please note that Frequency Transformation Blocks used in [10], published in USENIX, don’t work along the T-F axis (i.e., only time axis).
CGAB provides attention to the T-F axes of a complex-valued T-F spectrogram by following two steps (see Fig. 6):
@@ IV-E Complex Global Attention Block (CGAB)
Step 1 - Reshaping along the T-F axes: The encoder’s output $E_{n}$ is decomposed into two tensors: one along the time axis and another along the frequency axis. Formally, $E_{n}$, which has a feature dimension of $C\times F\times T$, is given at the input of CGAB. At the first stage of reshaping, $E_{n}$ is parallelly reshaped into $C.T$ vectors with dimension $C\cdot T\times F$ and into $C.F$ vectors with dimension $C\cdot F\times T$. This reshaping is done using 2D complex convolution, CBN, and complex ReLU activation, followed by vector reshaping. In the second stage of reshaping, $C\cdot T\times F$ is reshaped into $1\times T\times F$ and $C\cdot F\times T$ is reshaped into $1\times F\times T$ using 1D complex convolution, CBN, complex ReLU activation, followed by vector reshaping. The tensors with dimension $1\times F\times T$ capture the global harmonic correlation along the frequency axis, and $1\times T\times F$ capture the global inter-phoneme correlation along the time axis. The captured features along the T-F axes and the original features from $E_{n}$ are point-wise multiplied together to generate a combined feature map with a dimension of $C\times T\times F$ and $C\times F\times T$ along T and F axes, respectively. This point-wise multiplication captures the inter-channel relationship between the encoder’s output and complex T-F axes.
Step 2 - Global attention along the T-F axes: It is possible to treat the spectrogram as a 2D image and learn the correlations between every two pixels in the 2D image. However, this is computationally too costly and is not realistic. On the other hand, ideally, we can use self-attention [29] to learn the attention map from two consecutive complex T-F spectrograms. But this might not be necessary. Because, on the time axis in each T-F spectrogram, when calculating SNR, the same set of parameters in the recursive relation is used, which suggests that temporal correlation is time-invariant among consecutive spectrograms. Moreover, harmonic correlations are independent in the consecutive spectrograms [30].
@@
Based on this understanding, we propose a self-attention technique along the T-F axes. Specifically, attention to frequency and time axes is implemented by two separate fully connected (FC) layers. Along the time path, the input and output dimensions of FC layers are $C\times T\times F$. Along the frequency path, the input and output dimensions of FC layers are $C\times F\times T$. FC layer learns weights from complex T-F spectrograms and technically is different from the self-attention [29] operation. To capture interchannel relationships among the input $E_{n}$ and output of FC layers, concatenation occurs, followed by 2D complex convolutions, CBN, and complex ReLU activation. Finally, the learned weights from the T-F axes are concatenated together to form a unified tensor, which holds joint information on the T-F global correlations from each spectrogram.
We use only two CGABs - one after the 1st encoder, and another after the 7th encoder - to optimize the design.
IV-F Complex Multiresolution STFT Loss